OpenStudy (anonymous):

6th grade Math: write the phrase as an expression: twice a number z. My son didn't know if they meant 2 times z, or z times itself twice. I'm pretty sure the answer is 2z, but not sure how to tell him why.

OpenStudy (alexandervonhumboldt2):

ok let z be a number. twice z. twice=2. thus z+z=?

OpenStudy (anonymous):

Thank you. I always thought math got cumbersome when taken out of context. I'd rather see them working with simple equations than partial ones.

OpenStudy (whpalmer4):

yes, "twice a number \(z\)" is the same as \(2z\) "a number \(z\) multiplied by itself" or "a number \(z\) times itself" is \(z*z = z^2\) "a number \(z\) times itself twice" is actually \(z*z*z=z^3\) but probably someone saying that means \(z*z=z^2\)